# Finite presentability of outer automorphism groups of large-type Artin groups

### v1

@@ 4.2 Finitely Presented Stabilisers
It remains to study the kernel of $\rho$. In general this kernel is generated by “bitwists”, as defined in [Lev05].
@@
The fact that on each factor the bitwists act like inner automorphisms means $[D_{z,z^{\prime}}]\in\ker(\rho)\leq\mathrm{Out}^{T}(G)$ for each bitwist, where $T$ is the tree with respect to which the bitwist was defined, and $\rho$ is as above.
In our case, the group generated by the twists (or more accurately their images in $\mathrm{Out}(A)$), which we shall call $\mathcal{T}\leq\mathrm{Out}^{T}(A)$, is the entire kernel. Finite presentability of the kernel follows from Levitt’s explicit finite presentation of $\mathcal{T}$.
@@
Let $\mathrm{Out}^{T}(A)$ be a point stabiliser in $\mathcal{F}$. Then $\ker(\rho)$ is finitely presented.
@@
By [Lev05, Proposition 2.2], the bitwists generate $\ker(\rho)$. We will show that for each $[D_{z,z^{\prime}}]\in\mathrm{Out}^{T}(A)$, $[D_{z,z^{\prime}}]\in\mathcal{T}$, and thus $\ker(\rho)=\mathcal{T}$.
@@
We claim that for every $[D_{z,z^{\prime}}]\in\mathrm{Out}^{T}(A)$, the automorphism induced by $z$ (and thus $z^{\prime}$, as they induce the same automorphism) is inner. By Lemma 3.11 there are two cases for $G_{e}$. The first case is that $G_{e}$ is spherical parabolic dihedral (here Lemma 3.11 says this is true regardless of which isomorphism $A\cong A_{\Gamma}$ we pick, justifying our lack of choice of Artin presentation), but in this case $G_{e}\leq A$ is self-normalising by Lemma 2.9, so $z\in G_{e}$ by assumption that it normalises $G_{e}$, and of course it induces an inner automorphism. The second case is that $G_{e}=A_{s}$ where $s$ is a standard generator. However $s$ is not conjugate in $A$ to its inverse by Lemma 2.10, so in fact $z$ acts trivially. In either case, by the preceeding paragraph $[D_{z,z^{\prime}}]\in\mathcal{T}$. This completes the proof that $\ker(\rho)=\mathcal{T}$.
A presentation for $\mathcal{T}$ is given by [Lev05, Proposition 3.1]. In particular, $\mathcal{T}$ is written as a quotient of $\Pi_{e\in E}Z_{A_{o(e)}}(A_{e})$, where $E$ is the set of oriented edges of $T//A$ and $o(e)$ is the origin vertex. By Lemma 3.11, $A_{e}$ is always a spherical dihedral or cyclic parabolic subgroup of $A_{o(e)}$. Now $Z_{A_{o(e)}}(A_{e})$ is finitely presented: in the dihedral case this follows from Lemma 2.9 and the well known fact that the centre of a spherical dihedral Artin group is $\mathbb{Z}$; in the cyclic case it follows by Lemma 2.10. Since $|E|$ is finite, it follows that $\Pi_{e\in E}Z_{A_{o(e)}}(A_{e})$ is finitely presented. The quotient is by the product of the centres of the edge and vertex groups, of which there are finitely many and all are either $\mathbb{Z}$ or trivial. In particular, $\mathcal{T}$ is finitely presented, since it is the quotient of a finitely presented group by a finitely generated subgroup. ∎
@@
Let $S\subseteq\mathcal{F}$ be a simplex. Then $S$ has finitely presented stabiliser.
@@
Given that simplices in $\mathcal{F}$ correspond to chains of collapses, take $v\in S$ to be the top vertex in the partial order. Of course, $G_{S}\leq G_{v}$. Now, consider that there are only finitely many sequences of collapses of the $A$-tree corresponding to $v$, so $G_{v}S\subseteq\mathcal{F}$ is a finite subcomplex. It follows that $G_{S}\leq_{f.i.}G_{v}$, so it is sufficient to show that vertex stabilisers are finitely presented.
Now, the result follows from our work before: $G_{v}$ has a finite index subgroup which is finitely-presented-by-finitely-presented by Lemma 4.12 and Lemma 4.14, so is itself finitely presented. ∎

### v2

@@ 4.3 Type F stabilisers
It remains to study the kernel of $\rho_{T}$. In general this kernel is generated by “bitwists”, as defined in [Lev05].
@@
The fact that on each factor the bitwists act like inner automorphisms means $[D_{z,z^{\prime}}]\in\mathrm{Out}_{0,+}^{T}(G)\leq\mathrm{Out}^{T}(G)$ for each bitwist, where $T$ is the tree with respect to which the bitwist was defined.
In our case, the group generated by the twists (or more accurately their images in $\mathrm{Out}(A)$), which we shall call $\mathcal{T}\leq\mathrm{Out}^{T}(A)$, is the entirety of $\mathrm{Out}^{T}_{0,+}(A)$. This group is of type F due to Levitt’s explicit description of $\mathcal{T}$.
@@
Let $\mathrm{Out}^{T}(A)$ be a point stabiliser in $\mathcal{F}$. Then $\mathrm{Out}^{T}_{0,+}(A)$ is of type F.
@@
By [Lev05, Proposition 2.2], the bitwists generate $\mathrm{Out}^{T}_{0,+}(A)$. We will show that for each $[D_{z,z^{\prime}}]\in\mathrm{Out}^{T}(A)$, $[D_{z,z^{\prime}}]\in\mathcal{T}$, and thus $\mathrm{Out}^{T}_{0,+}(A)=\mathcal{T}$.
@@
We claim that for every $[D_{z,z^{\prime}}]\in\mathrm{Out}^{T}(A)$, the automorphism induced by $z$ (and thus $z^{\prime}$, as they induce the same automorphism) is inner. By Lemma 3.11 there are two cases for $G_{e}$. The first case is that $G_{e}$ is spherical parabolic dihedral (here Lemma 3.11 says this is true regardless of which isomorphism $A\cong A_{\Gamma}$ we pick, justifying our lack of choice of Artin presentation), but in this case $G_{e}\leq A$ is self-normalising by Lemma 2.10, so $z\in G_{e}$ by assumption that it normalises $G_{e}$, and of course it induces an inner automorphism. The second case is that $G_{e}=A_{s}$ where $s$ is a standard generator. However $s$ is not conjugate in $A$ to its inverse by Lemma 2.11, so in fact $z$ acts trivially. In either case, by the preceeding paragraph $[D_{z,z^{\prime}}]\in\mathcal{T}$. This completes the proof that $\mathrm{Out}^{T}_{0,+}(A)=\mathcal{T}$.
A presentation for $\mathcal{T}$ is given by [Lev05, Proposition 3.1]. In particular,
$\mathcal{T}=\Pi_{e\in E}Z_{A_{o(e)}}(A_{e})/(\Pi_{v\in V}Z(A_{v})\times\Pi_{e\in E^{+}}Z(A_{e})),$
where $E$ is the set of oriented edges of $T//A$ (and $E^{+}$ is a set of arbitrary representatives for each edge) and $o(e)$ is the origin vertex of such an edge. The centers of the edge and vertex stabilisers embed into the product of centralisers in the obvious diagonal way.
By Lemma 3.11, $A_{e}$ is always a spherical dihedral or cyclic parabolic subgroup of $A_{o(e)}$. If $A_{e}$ is dihedral then it follows from Lemma 2.10 that $Z_{A_{o(e)}}(A_{e})\cong\mathbb{Z}$ is the centre of $A_{e}$. If $A_{e}$ is cyclic then, by Lemma 2.11, $Z_{A_{o(e)}}(A_{e})\cong\mathbb{Z}\times F$, where $F$ is a finitely generated free group, and the central $\mathbb{Z}$ is exactly $Z(A_{e})=A_{e}$. In this case, if $A_{o(e)}$ is dihedral, then the $F$ direct factor is in fact cyclic, and generated by $Z(A_{o(e)})$.
So $\Pi_{e\in E}Z_{A_{o(e)}}(A_{e})$ is a direct product of $\mathbb{Z}$ and ×$\mathbb{Z}\times F$ factors. One easily checks from the preceding discussion that the quotient identifies some of the $\mathbb{Z}$ factors and free direct factors $F$ (when $F$ is cyclic). We conclude that $\mathcal{T}$ is a right angled Artin group and, in particular, is of type F. ∎
@@
Let $S\subseteq\mathcal{F}$ be a simplex. Then the stabiliser of $S$ under the action of $\mathrm{Out}_{0,+}(A)$ is of type F.
@@
Given that simplices in $\mathcal{F}$ correspond to chains of collapses, take $v\in S$ to be the top vertex in the partial order. Of course, $G_{S}\leq G_{v}$. Now, consider that there are only finitely many sequences of collapses of the $A$-tree corresponding to $v$, because collapses are equivariant and there are only finitely many orbits of edges. In particular, $G_{v}S\subseteq\mathcal{F}$ is a finite subcomplex. It follows that $G_{S}\leq_{f.i.}G_{v}$, and $G_{S}\cap\mathrm{Out}_{0,+}(A)\leq_{f.i.}G_{v}\cap\mathrm{Out}_{0,+}(A)$.
The result follows since $G_{v}\cap\mathrm{Out}_{0,+}(A)$ is of type F by Lemma 4.17, and being type F is closed under taking finite index subgroups. ∎